28. August 2001. U. Schoenwaelder; http://www.math.rwth-aachen.de/∼Ulrich.Schoenwaelder HB = Hochschulbibl. RWTH, HBZ = http://www.hbz-nrw.de/ (HBZ-CD-ROM Online), MB = Mathematikbibl., DB = Didaktikbibl. (Winter), FH = Bibl. Fachhochschule Aachen, FL = Fernleihe, IB Nr. Institutsbibliothek Nr., LB = HB–Lehrbuchsammlung, LS = HB–Lesesaal

Literatur zur Elementaren Zahlentheorie: Lehrbücher

[1] W. A. Adams and L. J. Goldstein. Introduction to Number Theory. Prentice–Hall, Inc., 1976. MB: 9094. Quadratic fields and forms. Chapters: 1 Introduction; 2 Divisibility and Primes; 3 Congruences; 4 The Law of Quadratic Reciprocity; 5 Arithmetic Functions (5.3 The Möbius Inversion Formula); 6 A Few Diophantine Equations; 7 The Gaussian Integers; 8 Arithmetic in Quadratic Fields; 9 Factorization Theory in Quadratic Fields; 10 Applications of the Factorization Theory to Diophantine Equations; 11 The Representation of Integers by Binary Quadratic Forms.

[2] A. Aigner. Zahlentheorie. de Gruyter, 1975. MB: 7882.

[3] P. S. Alexandroff, A. I. Markuschewitsch, and A. J. Chintschin. Enzyklopädie der Elementarmathematik, Band I: Arithmetik. Hochschulbücher für Mathematik. Deutscher Verlag der Wissenschaften, 1954. HB: 1 Bb 1054. S. 225–313: Die Elemente der Zahlentheorie (A. J. Chintschin): Kap. I: Teilbarkeit und Primzahlen; Kap. II: Die Kongruenzmethode; Kap. III: Der Euklidische Algorithmus und die Kettenbrüche; Kap. IV: Die Darstellung der Zahlen durch systematische Brüche und durch Kettenbrüche; Kap. V: Kettenbrüche und diophantische Approximationen; Kap. VI: Algebraische und transzendente Zahlen.

[4] W. S. Anglin. The Queen of Mathematics. An Introduction to Number Theory. Kluwer Texts in the Mathematical Sciences 8. Dordrecht: Kluwer, 1995. ISBN 0-7923-3287-3. HBZ. Tel. Rev.: Monthly 102:9 (1995), 852. 1. Propaedeutics. 2. Simple continued fractions. 3. Congruence. 4. x^{2}−Ry^{2} = C. 5. Classical construction problems. 6. The polygonal number theorem.

7. Analytic number theory.

[5] Andreas Bartholomé, Josef Rung, and Hans Kern. Zahlentheorie für Einsteiger. Vieweg-Mathematik für Schüler und Studenten. Vieweg, 1995, ^{2}1996. HBZ.

[6] J. W. Bruce, P. J. Giblin, and P. J. Rippon. Microcomputers and Mathematics. Cambridge Univ. Press, 1990. HB: Bm 2340. Euklidischer Algorithmus, Kettenbrüche, Fibonacci-Folge, Eulers Phi–Funktion, Legendre–Symbol, Primzahlen, Primzahltests, π, e (ist irrational: S. 278), Differentialgleichungssysteme.

[7] Peter Bundschuh. Einführung ind die Zahlentheorie. Hochschultext. Springer–Verlag, 1988, ^{3}1996. HB: Bf9021, Bf9021+3. ISBN 0-387-15305-5, 3-540-60920-2.

[8] S. A. Burr et al. The Unreasonable Effectiveness of Number Theory. Proceedings of Symposia in Applied Mathematics. AMS, 1993. Contents: M. R. Schroeder: The unreasonable effectiveness of number theory in physics, communication, and music; G. E. Andrews: The reasonable and unreasonable effectiveness of number theory in statistical mechanics; J. G. Lagarias: Number theory and dynamical systems; G. Marsaglia: The mathematics of random number generators; V. Pless: Cyclotomy and cyclic codes; M. D. McIlroy: Number theory in computer graphics.

[9] D. Burton. Elementary Number Theory. [New to] McGraw–Hill, 1997. HBZ. Contents: Divisibility Theory in the Integers / Primes and Their Distribution / The Theory of Congruences / Fermat’s Theorem / Primitive Roots and Indices / The Quadratic Reciprocity Law / Perfect Numbers / The Fermat Conjecture / Representation of Integers as Sums of Squares / Fibonacci Numbers / Continued Fractions / Some Twentieth Century Developments.

[10] Robin Chapman. A guide to arithmetic, August 5 1994. alg/arithm.rjc.dvi.

[11] A. J. Chinčin. Drei Perlen der Zahlentheorie. Akademie–Verlag, 1951. MB: 515.

[12] Judita Cofman. Einblicke in die Geschichte der Mathematik: Aufgaben und Materialien für die Sekundarstufe I. Texte zur Didaktik der Mathematik. Spektrum Akad. Verlag, 1999. Teil I: Aufgaben mit Bemerkungen zur Geschichte der Mathematik, Teil II: Lösungen, Teil III: Zusätzliche Bemerkungen aus der Geschichte der Mathematik.

[13] J. H. Conway and R. K. Guy. The Book of Numbers. Springer–Verlag, 1996. MB: 17923.

[14] H. Davenport. The Higher Arithmetic. An Introduction to the Theory of Numbers. Hutchinson & Co. Ltd.; Cambridge Univ. Press, 1952, ^{6}1992. MB: 16 723. Inhalt: I Factorization and primes. II Congruences. III Quadratic residues. IV Continued fractions. V Sums of squares. VI Quadratic forms. VII Some Diophantine equations. VIII Computers and the theory of numbers (7. The RSA cryptographic method).

[15] D. M. Davis. The Nature and Power of Mathematics. Princeton Univ. Press, 1993. Number theory with its application to cryptography.

[16] K. Denecke and K. Todorov. Algebraische Grundlagen der Arithmetik. Berliner Studienreihe zur Mathematik, Band 4. Heldermann Verlag, 1994. MB: 17179.

[17] H. M. Edgar. A First Course in Number Theory. Wadsworth, 1988. p–adic Newton’s method for successive approximations. Tel. Rev.: Amer. Math. Monthly 99:8(1992), 805.

[18] D. E. Flath. Introduction to Number Theory. John Wiley & Sons, 1989. MB: 14665. Quadratische Formen.

[19] P. Fletcher and C. W. Patty. Foundations of Higher Mathematics. PWS–Kent, ^{2}1992. Tel. Rev.: Amer. Math. Monthly 100(1):1993, 96.

[20] H. Freund. Elemente der Zahlentheorie. Teubner, 1994. MB: 10406.

[21] G. Frey. Elementare Zahlentheorie. Vieweg–Studium 56, Grundkurs Mathematik. Vieweg, 1984. MB: 13930.

[22] Jay R. Goldman. The Queen of Mathematics: A Historical(ly Motivated) Guide to Number Theory. A K Peters, 1997. ISBN 1- 56881-006-7. MB: 18414. Rekl.: Notices AMS 44:9 (1997), 1094.

[23] E. Grosswald. Topics from the Theory of Numbers. Birkhäuser, ^{2}1984. MB: 12306. Part Two: Elementary Number Theory.

[24] G. H. Hardy and E. M. Wright. Einführung in die Zahlentheorie. R. Oldenbourg, 1958. HB: Bf1585. Kap. 3: Fareyreihen und ein Satz von Minkowski. Original: An introduction to the Theory of Numbers, Oxford Univ. Press, ^{3}1954.

[25] G. H. Hardy and E. M. Wright. The Theory of Numbers. Oxford Univ. Press, ^{4}1960.

[26] F. Ischebeck. Einladung zur Zahlentheorie. BI Wissenschaftsverlag, 1992. MB: 16320. S. 186–187: Namensliste mit Lebensdaten.

[27] T. H. Jackson. From Number Theory to Secret Codes. Adam Hilger, 1987. Continued fractions and rational approximants.

[28] T. H. Jackson. From Polynomials to Sums of Squares. Inst. of Physics Publishing, 1995.

[29] M. Jeger. Computer–Streifzüge. Eine Einführung in Zahlentheorie und Kombinatorik aus algorithmischer Sicht. Programm–Praxis 5. Birkhäuser, 1986. HB: 5 Za 7362. Kap. 5: n!, π, e. Review: Praxis der Mathematik 31:1 (1989), 62–63.

[30] N. Koblitz. A Course in Number Theory and Cryptography. GTM 114. Springer–Verlag, 1987. HB: 114 Bb 1244. Ch. I: Some topics in elementary number theory; §1: Time estimates for doing arithmetic, §2: Divisibility and the Euclidean algorithm, §3: Congruences (Fermat’s Little Theorem, Chinese Remainder Theorem), §4: Some applications to factoring. Ch. II: Finite fields and quadratic residues. Ch. III: Cryptography. Ch. IV: Public Key. Ch. V: Primality and Factoring. Ch. VI: Elliptic Curves (cryptosystems, factorization).

[31] A. Leutbecher. Zahlentheorie. Eine Einführung in die Algebra. Grundwissen Mathematik, Springer–Lehrbuch. Springer– Verlag, 1996. MB: 17917.

[32] W. J. LeVeque. Topics in Number Theory, Volume 1. Addison–Wesley mathematics series. Addison–Wesley, 1956. MB: 592 a.

[33] W. J. LeVeque. Topics in Number Theory, Volume 2. Addison–Wesley mathematics series. Addison–Wesley, 1956. MB: 592 b.

[34] W. J. LeVeque. Elementary Theory of Numbers. Addison–Wesley series in introductory mathematics. Addison–Wesley, 1962. MB: 1716.

[35] Malcolm E. Lines. Think of a Number. Ideas, concepts and problems which challenge the mind and baffle the experts. Adam Hilger (IOP Publishing Ltd), 1990. FL: UB Bielefeld 100/1663205+1. ISBN 0-85274-183-9. Published also under the title: Numbers at work and at play.

[36] Bill McClung. Exploring numbers. http://archives.math.utk.edu/interactive.texts/projectpages/ en.html. Mit Mathematica.

[37] T. Nagell. Introduction to Number Theory. Wiley, 1951, ^{2}1981. MB: 4818. ISBN 0-8284-0163-2.

[38] I. Niven and H. S. Zuckerman. Einführung in die Zahlentheorie. BI, 1976. MB: 8911ab.

[39] I. Niven, H. S. Zuckerman, and H. L. Montgomery. An Introduction of the Theory of Numbers. Wiley, 1991.

[40] O. Ore. Invitation to Number Theory. MAA, 1967. Gentle introduction to elementary number theory.

[41] H.-H. Ostmann and H. Liermann. Zahlentheorie. In Behnke et al., editor, Grundzüge der Mathematik, number Bd. I: Grundlagen der Mathematik. Arithmetik und Algebra, chapter Teil B, Kapitel 6, pages 368–425. Göttingen: Vandenhoeck-Ruprecht, 1966. MB: 3947a. Vgl. [42].

[42] H.-H. Ostmann and H. Liermann. Theory of numbers. In H. Behnke et al., editor, Fundamentals of Mathematics, number Volume I: Foundations of Mathematics. The Real Number System and Algebra, chapter Part B, Chapter 6, pages 355–408. Cambridge, MA: MIT Press, 1974. ISBN 0-262-02048-3. MB: 8207a. Vgl. [41].

[43] F. Padberg. Didaktik der elementaren Zahlentheorie. smd. Herder, 1981. MB: 15714. Rev.: PM 25:12 (1983), 382.

[44] H. Rademacher. Lectures on Elementary Number Theory. Blaisdell Publ. Co., 1964. HB: Bf52.

[45] H. Rademacher. Lectures on Elementary Number Theory. R. E. Krieger Publ., 1977.

[46] Don Redmond. Number Theory - An Introduction. Pure and Appl. Math.: A Series of Monographs and Textbooks 201. Dekker, 1996. ISBN 0-8247-9696-9. HBZ.

[47] R. Remmert and P. Ullrich. Elementare Zahlentheorie. Birkhäuser, ^{2}1995. HB: Bf915+2 LB, B01303 Aufsicht LS. ISBN 3-7643-5197-7. Primfaktorzerlegung, ggT, Dezimalsystem, Kongruenzen, primitive Wurzeln, Reziprozitätsgesetz für quadratische Reste. Satz von Fermat-Euler in der Kryptographie.

[48] K. H. Rosen. Elementary Number Theory and Its Applications. Addison–Wesley, 1984. MB: 12714. Lamé’s theorem: S. 62. Euklidischer Algorithmus.

[49] H. Scheid. Zahlentheorie. BI, 1991. MB: 15713.

[50] Friedrich Schwarz. Einführung in die Elementare Zahlentheorie. MuPAD lectures. Teubner, 1998. HB: Bb5018. ISBN 3-519-02193-5. Mit CD-ROM für interaktive Version des Buches mit MuPad Light für Windows 95.

[51] J. Schwermer. Über Reziprozitätsgesetze in der Zahlentheorie. In H. Knörrer, C.-G. Schmidt, J. Schwermer, and P. Slodowy, editors, Arithmetik und Geometrie. Vier Vorlesungen, Mathematische Miniaturen 3, pages 29–69. Birkhäuser, 1986. MB: 13286.

[52] H. N. Shapiro. Introduction to the Theory of Numbers. Wiley, 1983. MB: 11800.

[53] H. Siemon. Anwendungen der elementaren Gruppentheorie in Zahlentheorie und Kombinatorik, chapter 2.8: Die Modulgruppe und ihre Anwendung auf binäre quadratische Formen, pages 83–102. Klett Studienbücher Mathematik. Ernst Klett, 1981. HB: Bf 7343. MB: 12232.

[54] W. Sierpiński. Elementary Theory of Numbers. North–Holland, 1988. MB: 14176. Umfassend. Ch. VIII: Continued fractions.

[55] H. Stark. An Introduction to Number Theory. Chicago: Markham, 1970.

[56] James T. Tattersall. Elementary Number Theory in Nine Chapters. Cambridge Univ. Press, 1999.

[57] B. A. Venkov. Elementary Number Theory. Wolters–Noordhoff, Groningen, 1970. MB: 5968.

[58] I. M. Winogradow. Elemente der Zahlentheorie. Hochschulbücher für Mathematik 22. VEB Deutscher Verlag der Wissenschaften, 1955.

[59] Song Y. Yan. Number Theory for Computing. Springer, 2000. MB: 18905. ISBN 3-540-65472-0. Elementary Number Theory: Divisibility; Diophantine Equations; Arithmetical Functions; Congruences; Arithmetic of Elliptic Curves. Algorithmic Number Theory: Primality Testing; Integer Factorization; Discrete Logarithms; Quantum Algorithmic Number Theory. Applied Number Theory: Computer Systems Design; Cryptology and Information Security.